# More Than Memory Savings: Zeroth-Order Optimization Mitigates Forgetting in Continual Learning

### v1

@@ 5 Experiments
LAE Adapter[15] 88.05 82.31 8.35 66.28 55.78 9.93 74.55 68.03 15.17
L2P[48] 85.05 78.59 9.09 69.60 62.35 6.12 76.39 70.30 12.22
CODA-Prompt[40] 87.58 80.60 7.04 64.54 55.23 5.30 78.84 73.29 10.20
APER Adapter[53] 87.90 81.68 6.01 71.52 63.42 7.48 71.24 64.93 6.72
EASE[54] 91.13 85.45 7.05 78.78 71.93 7.54 75.91 70.24 8.54
InfLoRA[25] 89.41 82.89 6.95 73.33 64.52 7.95 79.70 74.05 7.43
@@ 5.1 Experimental Details
We evaluate the methods on three widely-used benchmarks in CIL works: CIFAR100 [23], ImageNet-R [37] and DomainNet [33]. Following the settings in prior work [35, 40], we conduct the experiments using two settings: 5-class and 10-class per task for CIFAR100 and ImageNet-R, 69-class per task for DomainNet.
Evaluation Metric: Across all experiments, we report three key metrics following [44]: (1) average accuracy (avg) across all tasks at the end of training, (2) last-task accuracy (last) on the latest task, and (3) catastrophic forgetting (fgt), defined as the average drop in accuracy of old tasks from their best observed accuracy to their final accuracy [9]. Formal definitions are provided in Appendix A.
Training Details: All methods use a ViT-B/16 [8] backbone pre-trained on ImageNet-21K [37]. Following [15], we insert a LoRA-like adapter module with rank $r=5$ in parallel to the MLP layers in each of the first five transformer blocks. For ZO optimization, we set the perturbation magnitude to $\varepsilon=0.001$ and the query budget to $Q=4$ with L2 clipping at 1.0 for the ZO updates. In ZO-FC, the classifier is optimized with SGD (LR=0.01, cosine decay), and the adapter with ZO-SGD (LR=0.01, constant). Note that, we adopt the official implementations and follow the hyperparameter settings recommended in their original papers and the LAMDA-PILOT benchmark [44]. And the compared methods are conducted by using same checkpoints of the pre-trained model and task orders to ensure a fair comparison. The other related training configurations can be found in Appendix A.
@@ 5.3 Analysis and Ablation Study
ZO-FC finds flatter shared region to mitigate forgetting: To validate our analysis in Section 4, which shows that ZO-FC guides the model toward wider and flatter minima that mitigate forgetting, we quantify loss-landscape flatness during CL. While the curvature of the loss landscape can be characterized by Hessian eigenvalues, it is computationally expensive to compute them directly. We adopt an approach inspired by Sharpness-Aware Minimization (SAM) [12] as a more efficient alternative, which perturbs model parameters within a small neighborhood and measures the induced loss change. Concretely, for model parameters ($\theta$) and the old-task loss ($L_{\text{old}}$), we define the flatness metric as:
@@ 6 Conclusion
Appendix
@@ C. Proof Assumptions for ZO SPSA searches flat minima
Under the standard assumptions that $L$ is convex, three-times continuously differentiable, and its third derivatives are bounded (the usual $L_{3}$-smoothness” condition), Zhang et al. [51] proved that in $T=\tilde{\mathcal{O}}(d^{4}/\varepsilon^{2})$ iterations: SPSA drives the $\varepsilon$-smoothed loss $\tilde{L}_{\varepsilon}(\theta)$ to within $\varepsilon$ of its global minimum:
@@
where $\theta_{T}$ is the final parameters after $T$ SPSA steps. By the definition of $\tilde{L}_{\varepsilon}$, Zhang et al. [51] also showed
@@
Standard GD optimizes the raw loss $f$, which does not guarantee bias toward flatter minima. Empirically, SGD sometimes finds wide valleys due to noise, but this is incidental and depends on schedule and noise scale. Explicit methods such as Entropy-SGD [3], Sharpness-Aware Minimization (SAM) [12] bias GD trajectories into wide valleys by injecting perturbations into the update rule. In contrast, two-point ZO always optimizes the smoothed loss $\tilde{f}$, and theory shows it minimizes to within $\epsilon$ of the smallest-Hessian-trace solution. This explains why our ZO-FC retains low forgetting across various streams: it consistently biases the shared adapter toward flat representations, without extra regularization.
@@ F. Data Augmentation and Effects
All image augmentations are applied in the dataloader before the model forward. No augmentation is performed inside the forward pass. This pipeline is identical for FO, ZO, and ZO-FC. Stochastic layers in the backbone (e.g., dropout) are disabled during ZO updates. The detailed augmentations are shown below, following configurations in [44] and [25]:
@@ G. Further studies on Prompt-based Methods
We conduct further experiments on L2P and CODA-Prompt with more updates (100 epochs on CIFAR100, 200 on ImageNet-R). Despite the longer schedule, the two prompt based methods fail to converge under ZO optimization and their accuracy results remain far below adapter-based ZO training and ZO-FC. This matches recent studies [50] that prompt-based methods are harder to converge under ZO SPSA because optimizing prompt selection with penalties and contrastive alignment loss requires precise gradients.
@@ H. Memory and Computation Cost Analysis
Computation Cost Measurement: We use ViT-B/16 at $224^{2}$ with batch size B=48B{=}48. The measured forward FLOPs per image (fvcore) is $\mathbf{1.768\times 10^{10}}$, so per-batch forward FLOPs is $\mathbf{8.484\times 10^{11}}$. FO total is estimated via the common rule $\text{FLOPs}_{\text{FO,total}}\!\approx\!3\times\text{FLOPs}_{\text{forward}}$, yielding $\mathbf{2.545\times 10^{12}}$ per batch. ZO performs (2q+1)(2q{+}1) forward evaluations with no backward; at q=1q{=}1 this matches FO’s total, and at q=4q{=}4 it is $9\times$ the forward-only batch cost, i.e., $\mathbf{7.636\times 10^{12}}$. ZO-FC uses the same ZO forwards and adds a classifier-only backward during early epochs. Given our small $C$ and a frozen backbone, this extra term ($\alpha_{\text{FC}}\,B\,F_{\text{cls}}$) is orders of magnitude smaller than a ViT forward, so totals numerically coincide with ZO at the same $q$ (both early and late). This explicitly reports and quantifies the additional forward passes required by ZO and how training cost scales with $q$. In practice, parallelizing the $2q$ ZO forwards [14] can reduce wall-clock time without changing FLOPs and is compatible with ZO-FC for on-device continual learning.

### v2

@@ 5 Experiments
LAE Adapter[gao2023lae] 88.05 82.31 8.35 66.28 55.78 9.93 74.55 68.03 15.17
L2P[wang2022l2p] 85.05 78.59 9.09 69.60 62.35 6.12 76.39 70.30 12.22
CODA-Prompt[smith2023codap] 87.58 80.60 7.04 64.54 55.23 5.30 78.84 73.29 10.20
APER Adapter[zhou2024aper] 87.90 81.68 6.01 71.52 63.42 7.48 71.24 64.93 6.72
EASE[zhou2024ease] 91.13 85.45 7.05 78.78 71.93 7.54 75.91 70.24 8.54
InfLoRA[liang2024inflora] 89.41 82.89 6.95 73.33 64.52 7.95 79.70 74.05 7.43
@@ 5.1 Experimental Details
We evaluate the methods on three widely-used benchmarks in CIL works: CIFAR100 [krizhevsky2009learning], ImageNet-R [russakovsky2015imagenet] and DomainNet [peng2019moment]. Following the settings in prior work [rebuffi2017icarl, smith2023codap], we conduct the experiments using two settings: 5-class and 10-class per task for CIFAR100 and ImageNet-R, 69-class per task for DomainNet. We fix the class/task order using seed 1993 for all methods and datasets.
Evaluation Metric: Across all experiments, we report three key metrics following [sun2025pilot]: (1) average accuracy (avg) across all tasks at the end of training, (2) last-task accuracy (last) on the latest task, and (3) catastrophic forgetting (fgt), defined as the average drop in accuracy of old tasks from their best observed accuracy to their final accuracy [díazrodríguez2018dontforgetforgettingnew]. Formal definitions are provided in supplementary material.
Training Details: All methods use a ViT-B/16 [dosovitskiy2021an] backbone pre-trained on ImageNet-21K [russakovsky2015imagenet]. Following [gao2023lae], we insert a LoRA-like adapter module with rank $r=5$ in parallel to the MLP layers in each of the first five transformer blocks. For ZO optimization, we set the perturbation magnitude to $\varepsilon=0.001$ and the query budget to $Q=4$ with L2 clipping at 1.0 for the ZO updates. In ZO-FC, the classifier is optimized with SGD (LR=0.01, cosine decay), and the adapter with ZO-SGD (LR=0.01, constant). Note that, we adopt the official implementations and follow the hyperparameter settings recommended in their original papers and the LAMDA-PILOT benchmark [sun2025pilot]. And the compared methods are conducted by using same checkpoints of the pre-trained model and task orders to ensure a fair comparison. The other related training configurations can be found in supplementary material.
@@ 5.3 Analysis and Ablation Study
ZO-FC finds flatter shared region to mitigate forgetting: To validate our analysis in Section 4, which shows that ZO-FC guides the model toward wider and flatter minima that mitigate forgetting, we quantify loss-landscape flatness during CL. While the curvature of the loss landscape can be characterized by Hessian eigenvalues, it is computationally expensive to compute them directly. We adopt an approach inspired by Sharpness-Aware Minimization (SAM) [foret2021sharpnessaware] as a more efficient alternative, which perturbs model parameters within a small neighborhood and measures the induced loss change. Concretely, for model parameters ($\theta$) and the old-task loss ($L_{\text{old}}$), we define the flatness metric as:
@@ 6 Conclusion
Acknowledgments
This work was supported in part by the U.S. National Science Foundation under Grant No.2348376.
Supplementary Material
@@ C. Proof Assumptions for ZO SPSA searches flat minima
Under the standard assumptions that $L$ is convex, three-times continuously differentiable, and its third derivatives are bounded (the usual $L_{3}$-smoothness” condition), zhang2025zofindsflat proved that in $T=\tilde{\mathcal{O}}(d^{4}/\varepsilon^{2})$ iterations: SPSA drives the $\varepsilon$-smoothed loss $\tilde{L}_{\varepsilon}(\theta)$ to within $\varepsilon$ of its global minimum:
@@
where $\theta_{T}$ is the final parameters after $T$ SPSA steps. By the definition of $\tilde{L}_{\varepsilon}$, zhang2025zofindsflat also showed
@@
Standard GD optimizes the raw loss $f$, which does not guarantee bias toward flatter minima. Empirically, SGD sometimes finds wide valleys due to noise, but this is incidental and depends on schedule and noise scale. Explicit methods such as Entropy-SGD [chaudhari2019entropy], Sharpness-Aware Minimization (SAM) [foret2021sharpnessaware] bias GD trajectories into wide valleys by injecting perturbations into the update rule. In contrast, two-point ZO always optimizes the smoothed loss $\tilde{f}$, and theory shows it minimizes to within $\epsilon$ of the smallest-Hessian-trace solution. This explains why our ZO-FC retains low forgetting across various streams: it consistently biases the shared adapter toward flat representations, without extra regularization.
@@ F. Data Augmentation and Effects
All image augmentations are applied in the dataloader before the model forward. No augmentation is performed inside the forward pass. This pipeline is identical for FO, ZO, and ZO-FC. Stochastic layers in the backbone (e.g., dropout) are disabled during ZO updates. The detailed augmentations are shown below, following configurations in [sun2025pilot] and [liang2024inflora]:
@@ G. Further studies on Prompt-based Methods
We conduct further experiments on L2P and CODA-Prompt with more updates (100 epochs on CIFAR100, 200 on ImageNet-R). Despite the longer schedule, the two prompt based methods fail to converge under ZO optimization and their accuracy results remain far below adapter-based ZO training and ZO-FC. This matches recent studies [yang2024adazeta] that prompt-based methods are harder to converge under ZO SPSA because optimizing prompt selection with penalties and contrastive alignment loss requires precise gradients.
@@ H. Memory and Computation Cost Analysis
Computation Cost Measurement: We use ViT-B/16 at $224^{2}$ with batch size B=48B{=}48. The measured forward FLOPs per image (fvcore) is $\mathbf{1.768\times 10^{10}}$, so per-batch forward FLOPs is $\mathbf{8.484\times 10^{11}}$. FO total is estimated via the common rule $\text{FLOPs}_{\text{FO,total}}\!\approx\!3\times\text{FLOPs}_{\text{forward}}$, yielding $\mathbf{2.545\times 10^{12}}$ per batch. ZO performs (2q+1)(2q{+}1) forward evaluations with no backward; at q=1q{=}1 this matches FO’s total, and at q=4q{=}4 it is $9\times$ the forward-only batch cost, i.e., $\mathbf{7.636\times 10^{12}}$. ZO-FC uses the same ZO forwards and adds a classifier-only backward during early epochs. Given our small $C$ and a frozen backbone, this extra term ($\alpha_{\text{FC}}\,B\,F_{\text{cls}}$) is orders of magnitude smaller than a ViT forward, so totals numerically coincide with ZO at the same $q$ (both early and late). This explicitly reports and quantifies the additional forward passes required by ZO and how training cost scales with $q$. In practice, parallelizing the $2q$ ZO forwards [gao2024enabling] can reduce wall-clock time without changing FLOPs and is compatible with ZO-FC for on-device continual learning.